### v1

@@ Abstract
Neural networks with piecewise linear activation functions, such as rectified linear units (ReLU) or maxout, are among the most fundamental models in modern machine learning. We make a step towards proving lower bounds on the size of such neural networks by linking their representative capabilities to the notion of the extension complexity $\operatorname{xc}(P)$ of a polytope $P$, a well-studied quantity in combinatorial optimization and polyhedral geometry. To this end, we propose the notion of virtual extension complexity $\operatorname{vxc}(P)=\min\{\operatorname{xc}(Q)+\operatorname{xc}(R)\mid P+Q=R\}$. This generalizes $\operatorname{xc}(P)$ and describes the number of inequalities needed to represent the linear optimization problem over $P$ as a difference of two linear programs. We prove that $\operatorname{vxc}(P)$ is a lower bound on the size of a neural network that optimizes over $P$. While it remains open to derive strong lower bounds on virtual extension complexity, we show that powerful results on the ordinary extension complexity can be converted into lower bounds for monotone neural networks, that is, neural networks with only nonnegative weights. Furthermore, we show that one can efficiently optimize over a polytope $P$ using a small virtual extended formulation. We therefore believe that virtual extension complexity deserves to be studied independently from neural networks, just like the ordinary extension complexity. As a first step in this direction, we derive an example showing that extension complexity can go down under Minkowski sum.
Keywords:
Neural Networks Polyhedral Combinatorics Linear Optimization Minkowski Sum Extension Complexity.
1 Introduction
@@
While machine learning is nowadays driven by rapid empirical progress, our work focuses on the theoretical understanding of the underlying models and methods. A crucial building block are feedforward neural networks. These are directed, acyclic graphs in which each vertex (neuron) defines a simple computation, usually a linear transformation composed with a scalar-valued, continuous and piecewise linear (CPWL) activation function. While a standard choice for the activation function is the rectified linear unit (ReLU) $x\mapsto\max\{0,x\}$, in this paper we focus on the more general maxout networks. These allow to compute the maximum of constantly many linear functions at each neuron. As a result, the entire network computes a (potentially complex) CPWL function. One of the big challenges in the theoretical analysis of neural networks is to understand how many neurons one requires to exactly or approximately represent a given (CPWL) function. To the best of our knowledge, it is an open question whether there exists a family of CPWL functions, which we can evaluate in polynomial time, but which cannot be represented by polynomial-size neural networks.
@@
The piecewise linear nature of the studied networks suggests to tackle such questions by means of polyhedral geometry, see, e.g., the recent survey [25]. In fact, a similar problem to the question above used to be open for a long time in the context of linear programming, until Rothvoß [37] resolved it affirmatively: does there exist a polytope $P$ over which we can optimize in polynomial time, but any linear programming formulation must have exponential size? This question can be formalized with the notion of extension complexity $\operatorname{xc}(P)$, which describes the minimal number of facets of any polytope $Q$ that projects onto $P$. In this case, we call $Q$ an extended formulation of $P$. Rothvoß [37] proved that the matching polytope has exponential extension complexity even though Edmonds’ algorithm [11] can be used to find the maximum weight matching or minimum weight perfect matching in polynomial time. Additionally, it was also proven that a couple of polytopes associated with NP-hard optimization problems like the travelling salesperson problem have exponential extension complexity [14].
@@
There is a direct translation between polytopes (as feasible sets of linear programs) and CPWL functions (represented by neural networks) through the notion of the support function $f_{P}(c)=\max_{x\in P}c^{\top}x$ of a polytope $P$. This (convex) CPWL function, which has one linear region for each vertex of $P$, uniquely determines $P$ via convex duality. Computing $f_{P}(c)$ means determining the objective value when optimizing over $P$ in $c$-direction. Each CPWL function $f_{P}$ can be represented by a neural network [1], though the required number of neurons can be large. To quantify this, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a maxout neural network.
1.1 Our Contributions
@@
The aim of this paper is to connect the world of extended formulations with the study of neural networks. A “dream result” in this direction would be to bound $\operatorname{xc}(P)$ polynomially in $\operatorname{nnc}(P)$. Then the breakthrough results on extension complexity would directly imply strong lower bounds on the size of neural networks. It turns out, however, that there is one feature of neural networks that seems to make the “dream result” difficult to obtain: namely the ability to use subtraction. To remedy this, we propose the notion of virtual extension complexity
@@
where $P+Q=\{p+q\mid p\in P,q\in Q\}$ is the Minkowski sum. In this definition, $P$ is a (formal) Minkowski difference of two polytopes $Q$ and $R$. Note that this is not the same as $R+(-1)\cdot Q$ but rather the inverse operation of Minkowski addition. The name virtual extension complexity is derived from virtual polytopes [33], a framework for the algebraic study of formal Minkowski differences of polytopes. Observe that $\operatorname{vxc}(P)\leq\operatorname{xc}(P)$ because we can always choose $Q=\{0\}$ with $\operatorname{xc}(\{0\})=0$. In that sense, virtual extension complexity is really a strengthening of the ordinary extension complexity.
In Section 3, we deduce that $\operatorname{vxc}(P)$ is indeed a lower bound for $\operatorname{nnc}(P)$, up to a constant factor. This leaves the open question to find ways to lower-bound $\operatorname{vxc}(P)$ in order to achieve the original goal to lower-bound $\operatorname{nnc}(P)$.
A consequence of the discussion above is that, if one removes the ability to subtract within neural networks, we do indeed obtain our “dream result”, namely lower bounds through $\operatorname{xc}(P)$. The resulting model is the one of monotone neural networks, which are neural networks with only nonnegative weights. In analogy to $\operatorname{nnc}(P)$, we define the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a monotone maxout neural network. In Section 4, we show how lower bounds on $\operatorname{xc}(P)$ imply lower bounds on exact and approximate representations of monotone neural networks.
Studying monotone neural networks is justified both from a theoretical and a practical perspective. From the theoretical perspective, it is a natural approach in complexity theory to prove lower bounds first for monotone models of computation, to first circumvent some additional challenges of the general case; see, e.g., [42]. From the practical perspective, it is always a good idea to incorporate prior knowledge about a problem into a machine learning algorithm to solve that problem. If one knows that the target function should be a monotone function, then it might be a good idea to constrain each unit of the neural network to compute a monotone function—this exactly corresponds to nonnegative weights. See [29] and the references therein for recent studies of monotone neural networks in the machine learning community.
Summarizing, in this paper, we discuss four ways to represent a polytope or its support function through (virtual) extended formulations and (monotone) neural networks. Figure 1 shows what we know about how the associated complexity measures $\operatorname{xc}(P)$, $\operatorname{vxc}(P)$, $\operatorname{nnc}(P)$, and $\operatorname{mnnc}(P)$ relate to each other.
@@
Figure 1: Relations between complexity measures for a polytope $P$. A directed arc means that the tail can be polynomially bounded by the head. It remains an open question whether $\operatorname{xc}(P)$ and $\operatorname{nnc}(P)$ can be related this way.
We conclude with some insights for $\operatorname{vxc}(P)$ and its relation to $\operatorname{xc}(P)$. Observe that $P+Q=R$ is equivalent to $f_{P}=f_{R}-f_{Q}$ pointwise. Therefore, intuitively, in order to optimize over $P$, one only needs to optimize over $R$ and $Q$ and subtract the results. We make this intuition formal in Section 5.1, implying that small extended formulations for $Q$ and $R$ are sufficient to optimize efficiently over $P$. Furthermore, in Section 5.2, we provide a class of examples with $P+Q=R$ demonstrating that $\operatorname{xc}(R)$ can be much smaller than $\operatorname{xc}(P)$. This gives important insights on how extension complexity behaves under Minkowski sum and implies that we really need to look at both $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$ in order to lower-bound $\operatorname{vxc}(P)$.
1.2 Further Related Work
@@
In this paper we aim to prove lower bounds on neural networks representing combinatorial optimization problems. Complementing upper bounds can be found in [22] for minimum spanning trees and maximum flows. Furthermore, in [23], similar upper bounds for the knapsack problem are given, even though they are different in flavor because some integrality assumptions are made.
@@
Concerning the general expressivity of (piecewise linear) neural networks, the celebrated universal approximation theorems state that a single layer of neurons is sufficient to approximate any continuous function on a bounded domain; see [10] for the original version for sigmoid activation functions and [27] for a version that encompasses ReLU. However, such shallow neural networks usually require a large number of neurons. A sequence of results demonstrates that deeper networks sometimes require exponentially fewer neurons to represent the same functions; see, e.g., [1, 12]. In terms of exact representation, it is known that a function can be represented if and only if it is CPWL [1], and it is still an open question whether constant depth is sufficient to do so [21, 20]. Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts [41]. Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently [5, 40].
@@
We would like to emphasize that we view neural networks as a model of real-valued computation, as opposed to binary models of computation like Boolean circuits and Turing machines. In fact, if one restricts the inputs of a neural network to be binary, it is not too difficult to simulate AND-, OR-, and NOT-gates [32]. Thus, in such a binary model, every problem in P can be solved with polynomial-size neural networks. However, such networks would usually be very sensitive to single bits in the input and therefore cannot naturally be transformed into exact or approximate neural networks in the real-valued model. See also the discussion in [22] for more details.
The more useful connection to circuit complexity is through arithmetic [38], and in particular tropical circuits [26], which are also real-valued models of computation. Again we refer to [22] for a more detailed discussion of how these models relate to neural networks.
@@
In fact, the extension complexity has been related before to Boolean and arithmetic circuits, see [13, 24]. This is also related to Goemans’ proof that the permutahedron has extension complexity $\mathcal{O}(n\log n)$ [16], as this goes via sorting networks, which can be seen as a very specific version of a piecewise-linear arithmetic circuit.
2 Preliminaries
@@
Each of $V$- and $H$-representation of a polytope can be exponentially smaller than the other one. For example, the $d$-dimensional cube $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{\infty}\leq 1\}$ has $2^{d}$ vertices but only $2d$ facets. On the other hand, the $d$-dimensional cross-polytope $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{1}\leq 1\}$ has only $2d$ vertices, but $2^{d}$ facets. Standard references for polytope theory are [19, 45].
2.0.1 Extended Formulations.
Sometimes, a more compact way of representing a polyhedron $P$ can be obtained by representing it as the projection of a higher-dimensional polyhedron $Q\subseteq\mathbb{R}^{e}$ with $e\geq d$. Such a polytope $Q$ with $\pi(Q)=P$ for an affine projection $\pi$ is called an extended formulation of $P\subseteq\mathbb{R}^{d}$. The extension complexity of $P$ is defined as $\operatorname{xc}(P)=\min\{h(Q)\mid\pi(Q)=P\}$. The extension complexity is upper bounded by $\min\{v(P),h(P)\}$, but can be exponentially smaller. For example, the spanning tree polytope of a graph with $n$ vertices has extension complexity $\mathcal{O}(n^{3})$ [28], but the polytope itself has exponentially many vertices and facets in $n$. See [9] for a (not very recent) survey on extended formulations.
2.0.2 Support Functions.
@@
The inverse map that maps a support function $f_{P}$ to its unique associated polytope $P$ is well-studied in tropical geometry. In fact, borrowing the name from (tropical) polynomials, we call $P$ the Newton polytope of $f_{P}$. The connection between CPWL functions and their Newton polytopes has previously been used in the study of neural networks and their expressivity [44, 31, 20, 21, 6, 30].
2.0.3 Neural Networks.
For an introduction to neural networks, we refer to [17]. In this paper, we focus on so-called rank-$k$ maxout neural networks for some natural number $k\geq 2$. They generalize the well-known neural networks with rectified linear unit (ReLU) activations. Such a neural network is based on a directed, acyclic graph with node set $V$ and arcs $A\subseteq V\times V$, where the nodes are called neurons and act as computational units. The $d\geq 1$ neurons with in-degree zero are called input neurons, all other $s\geq 1$ neurons are maxout units. We assume that among the maxout units there exists exactly one neuron with out-degree zero, the output neuron. All neurons that are neither input nor output neurons are called hidden neurons. The size of the neural network is defined as the number $s$ of maxout units.
Each neuron $v$ in the neural network defines a function $z_{v}\colon\mathbb{R}^{d}\to\mathbb{R}$ as follows. Each input neuron $v$ is associated with one element $x_{i}$ of the input vector $x\in\mathbb{R}^{d}$ and simply outputs $z_{v}(x)=x_{i}$. Each maxout unit $v$ comes with a tuple of weights $w_{uv}^{i}\in\mathbb{R}$ for $i=1,\dots,k$ and all $u\in\delta^{\mathrm{in}}_{v}$, where $\delta^{\mathrm{in}}_{v}$ is the set of in-neighbors of $v$. They are naturally associated with $k$ weight vectors $w_{*v}^{i}=(w_{uv}^{i})_{u\in\delta^{\mathrm{in}}_{v}}$ of length $\lvert\delta^{\mathrm{in}}_{v}\rvert$ for $i=1,\dots,k$. The maxout neuron $v$ represents the following expression dependent on the outputs of its in-neighbors:
@@
Finally, the output of the network is defined to be the output $z_{v}$ for the output neuron $v$. Note that our neural network represents a scalar-valued function as we assumed that there is only a single output neuron. We call a maxout network monotone if all weights of all maxout neurons are nonnegative.
For a polytope $P$, we define the neural network complexity $\operatorname{nnc}(P)$ and the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum size of a (monotone) rank-$2$ maxout network representing $f_{P}$.
Remark 1 (biases).
In practice, the units usually also involve biases. In the following, we will mainly focus on support functions of polytopes, which are positively homogeneous. Therefore, we can omit the bias in the definition of neural networks without loss of generality, compare [21, Proposition 2.3]. With our definition of a maxout network, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ can be represented by such a network if and only if it is CPWL and positively homogeneous.
Remark 2 (maxout ranks and ReLU).
@@
3 Neural Networks are Virtual Extended Formulations
In this section we prove the following theorem stating that neural network sizes can be lower-bounded through virtual extension complexity. Here and in the following, we defer some of the proofs to Appendix 0.A.

### v2

@@ Abstract.
Neural networks with piecewise linear activation functions, such as rectified linear units (ReLU) or maxout, are among the most fundamental models in modern machine learning. We make a step towards proving lower bounds on the size of such neural networks by linking their representative capabilities to the notion of the extension complexity $\operatorname{xc}(P)$ of a polytope $P$. This is a well-studied quantity in combinatorial optimization and polyhedral geometry describing the number of inequalities needed to model $P$ as a linear program. We show that $\operatorname{xc}(P)$ is a lower bound on the size of any monotone or input-convex neural network that solves the linear optimization problem over $P$. This implies exponential lower bounds on such neural networks for a variety of problems, including the polynomially solvable maximum weight matching problem.
In an attempt to prove similar bounds also for general neural networks, we introduce the notion of virtual extension complexity $\operatorname{vxc}(P)$, which generalizes $\operatorname{xc}(P)$ and describes the number of inequalities needed to represent the linear optimization problem over $P$ as a difference of two linear programs. We prove that $\operatorname{vxc}(P)$ is a lower bound on the size of any neural network that optimizes over $P$. While it remains an open question to derive useful lower bounds on $\operatorname{vxc}(P)$, we argue that this quantity deserves to be studied independently from neural networks by proving that one can efficiently optimize over a polytope $P$ using a small virtual extended formulation.
1. Introduction
@@
A feedforward neural network is a directed, acyclic graph in which each vertex (neuron) defines a simple computation, usually a linear transformation composed with a scalar-valued, continuous and piecewise linear (CPWL) activation function. While a standard choice for the activation function is the rectified linear unit (ReLU) $x\mapsto\max\{0,x\}$, in this paper we focus on the more general maxout networks. These allow to compute the maximum of constantly many linear functions at each neuron. As a result, the entire network computes a (potentially complex) CPWL function. One of the big challenges in the theoretical analysis of neural networks is to understand how many neurons one requires to exactly or approximately represent a given (CPWL) function. To the best of our knowledge, it is an open question whether there exists a family of CPWL functions, which we can evaluate in polynomial time, but which cannot be represented by polynomial-size neural networks. [^1] Note that here we refer to exact real-valued and not binary computation, compare the related discussion in Section 1.2.
@@
The piecewise linear nature of the studied networks suggests to tackle such questions by means of polyhedral geometry, see, e.g., the recent survey by Huchette et al. (2023). In fact, a similar problem to the question above used to be open for a long time in the context of linear programming, until Rothvoß (2017) resolved it affirmatively: does there exist a polytope $P$ over which we can optimize in polynomial time, but any linear programming formulation must have exponential size? This question can be formalized with the notion of extension complexity $\operatorname{xc}(P)$, which describes the minimal number of facets of any polytope $Q$ that projects onto $P$. In this case, we call $Q$ an extended formulation of $P$. Rothvoß (2017) proved that the matching polytope has exponential extension complexity even though the algorithm by Edmonds (1965) can be used to find the maximum weight matching or minimum weight perfect matching in polynomial time. Additionally, Fiorini et al. (2015) proved that a couple of polytopes associated with NP-hard optimization problems like the traveling salesperson problem have exponential extension complexity. Rothvoß (2017) and Fiorini, Massar, Pokutta, Tiwary, and De Wolf (2015) received the Gödel Prize 2023 for their breakthrough results.
@@
There is a direct translation between polytopes (as feasible sets of linear programs) and CPWL functions (represented by neural networks) through the notion of the support function $f_{P}(c)=\max_{x\in P}c^{\top}x$ of a polytope $P$. This (convex) CPWL function, which has one linear region for each vertex of $P$, uniquely determines $P$ via convex duality. Computing $f_{P}(c)$ means determining the objective value when optimizing over $P$ in $c$-direction. Each CPWL function $f_{P}$ can be represented by a neural network (Arora et al., 2018), though the required number of neurons can be large. To quantify this, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a maxout neural network.
1.1. Our Contributions
@@
The aim of this paper is to connect the world of extended formulations with the study of neural networks. A “dream result” in this direction would be to bound $\operatorname{xc}(P)$ polynomially in $\operatorname{nnc}(P)$. Then the breakthrough results on extension complexity would directly imply strong lower bounds on the size of neural networks. It turns out, however, that there is one feature of neural networks that seems to make the “dream result” difficult to obtain or maybe even wrong: namely the ability to use subtraction. There are two natural ways to circumvent this difficulty: either prove the “dream result” for a weaker version of neural networks, or prove it for a stronger version of extension complexity. We accomplish both of these variants in our paper.
Monotone and Input-Convex Neural Networks
A consequence of the discussion above is that, if we remove the ability to subtract within neural networks, we do indeed obtain our “dream result”, namely lower bounds through $\operatorname{xc}(P)$. This leads to two different, but closely related neural network models: monotone (Mikulincer and Reichman, 2022) and input-convex neural networks (ICNNs; Amos et al. (2017)). The former only allows nonnegative weights, therefore enforcing each neuron and the entire network to represent a monotone and convex function. The latter allows negative weights only on outgoing connections from the input neurons, therefore also enforcing the represented function to be convex, but no longer monotone.
Obviously, every network that is monotone in the sense defined above is also input-convex. In Section 3, we show that for computing monotone functions with maxout activations, ICNNs are not more efficient than monotone networks. So, in this case, both models are equivalent. In analogy to $\operatorname{nnc}(P)$, we define the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a monotone network or an ICNN. [^2] Even though ICNNs are used more frequently in practice, we still prefer to name the complexity measure after monotone networks due to the significance of monotone models in the circuit complexity community. We then show that exponential lower bounds on $\operatorname{xc}(P)$ imply corresponding lower bounds on $\operatorname{mnnc}(P)$ and consequently on exact and approximate representations with monotone networks and ICNNs.
Studying these restricted types of neural networks is justified both from a theoretical and a practical perspective. From the theoretical perspective, it is a natural approach in complexity theory to prove lower bounds first for monotone models of computation, to circumvent some additional challenges of the general case; see, e.g., Valiant (1979). From the practical perspective, it is sometimes desirable or even necessary to incorporate prior knowledge about monotonicity and/or convexity of the target function into a machine learning model. For convexity, ICNNs have been used extensively for exactly this purpose after they were introduced by Amos et al. (2017); see, e.g., the literature overview in Gagneux et al. (2025). For monotonicity, we refer to Mikulincer and Reichman (2022) and the references therein for recent studies of monotone neural networks in the machine learning community.
Depending on whether similar lower bounds can also be established for general neural networks or not, our exponential lower bounds on monotone networks and ICNNs could indicate that these models might not be the best solution for enforcing convexity of a machine learning model. This aligns with some observations by Gagneux et al. (2025).
Virtual Extension Complexity.
In order to pave the way towards lower bounds for general neural networks, we propose the notion of virtual extension complexity
@@
where $P+Q=\{p+q\mid p\in P,q\in Q\}$ is the Minkowski sum. In this definition, $P$ is a (formal) Minkowski difference of two polytopes $Q$ and $R$. Note that this is not the same as $R+(-1)\cdot Q$ but rather the inverse operation of Minkowski addition. The name virtual extension complexity is derived from virtual polytopes (Panina and Streinu, 2015), a framework for the algebraic study of formal Minkowski differences of polytopes. Observe that $\operatorname{vxc}(P)\leq\operatorname{xc}(P)$ because we can always choose $Q=\{0\}$ with $\operatorname{xc}(\{0\})=0$. In that sense, virtual extension complexity is really a strengthening of the ordinary extension complexity.
In Section 4.1, we deduce that $\operatorname{vxc}(P)$ is indeed a lower bound for $\operatorname{nnc}(P)$, up to a constant factor. This leaves the open question to find ways to lower-bound $\operatorname{vxc}(P)$ in order to achieve the original goal to lower-bound $\operatorname{nnc}(P)$. To this end, it seems to be crucial to generally obtain a better understanding of the complexity measure $\operatorname{vxc}(P)$ and its relation to $\operatorname{xc}(P)$. Observe that $P+Q=R$ is equivalent to $f_{P}=f_{R}-f_{Q}$ pointwise. Therefore, intuitively, in order to optimize over $P$, one only needs to optimize over $R$ and $Q$ and subtract the results. We make this intuition formal in Section 4.2, implying that small extended formulations for $Q$ and $R$ are sufficient to optimize efficiently over $P$. Furthermore, in Section 4.3, we provide a class of examples with $P+Q=R$ demonstrating that $\operatorname{xc}(R)$ can be much smaller than $\operatorname{xc}(P)$. This gives important insights on how extension complexity behaves under Minkowski sum and implies that we really need to look at both $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$ in order to lower-bound $\operatorname{vxc}(P)$.
Overall, we now have four different ways to represent a polytope or its support function through (virtual) extended formulations and (monotone) neural networks. Figure 1 shows what we know about how the associated complexity measures $\operatorname{xc}(P)$, $\operatorname{vxc}(P)$, $\operatorname{nnc}(P)$, and $\operatorname{mnnc}(P)$ relate to each other.
@@
Figure 1. Relations between complexity measures for a polytope $P$. A directed arc means that the tail can be polynomially bounded by the head. It remains an open question whether $\operatorname{xc}(P)$ and $\operatorname{nnc}(P)$ can be related this way.
1.2. Further Related Work
@@
In this paper our lever to prove lower bounds on neural networks are combinatorial optimization problems. Complementing upper bounds were established by Hertrich and Sering (2024) for minimum spanning trees and maximum flows. Furthermore, Hertrich and Skutella (2023) proved similar upper bounds for the knapsack problem, even though they are different in flavor because some integrality assumptions are made.
@@
Concerning the general expressivity of (piecewise linear) neural networks, the celebrated universal approximation theorems state that a single layer of neurons is sufficient to approximate any continuous function on a bounded domain; see Cybenko (1989) for the original version for sigmoid activation functions and Leshno et al. (1993) for a version that encompasses ReLU. However, such shallow neural networks usually require a large number of neurons. A sequence of results demonstrates that deeper networks sometimes require exponentially fewer neurons to represent the same functions; see, e.g., Arora et al. (2018); Eldan and Shamir (2016); Telgarsky (2016). While these works contain exponential lower bounds on the size of neural networks, they are focused on shallow networks. In contrast, we aim to prove lower bounds regardless of the depth.
In terms of exact representation, it is known that a function can be represented if and only if it is CPWL (Arora et al., 2018), and it is still an open question whether constant depth is sufficient to do so (Hertrich et al., 2023; Haase et al., 2023). Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts (Valerdi, 2024). Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently (Brandenburg et al., 2024a; Tran and Wang, 2024).
@@
We would like to emphasize that we view neural networks as a model of real-valued computation, as opposed to binary models of computation like Boolean circuits and Turing machines. In fact, if one restricts the inputs of a neural network to be binary, it is not too difficult to simulate AND-, OR-, and NOT-gates (Mukherjee and Basu, 2017). Thus, in such a binary model, every problem in P can be solved with polynomial-size neural networks. However, such networks would usually be very sensitive to single bits in the input. This is undesirable for practical neural networks and makes it impossible to transform these constructions naturally into exact or approximate neural networks in the real-valued model, compare the discussion by Hertrich and Sering (2024). The more useful connection to circuit complexity is through arithmetic (Shpilka et al., 2010), and in particular tropical circuits (Jukna, 2023), which are also real-valued models of computation. Again we refer to Hertrich and Sering (2024) for a more detailed discussion.
@@
In fact, the extension complexity has been related before to Boolean and arithmetic circuits, see Fiorini et al. (2021); Hrubeš and Yehudayoff (2023). This is also related to the proof that the permutahedron has extension complexity $\mathcal{O}(n\log n)$ (Goemans, 2015), as this goes via sorting networks, which can be seen as a very specific version of a piecewise-linear arithmetic circuit.
2. Preliminaries
@@
Each of $V$- and $H$-representation of a polytope can be exponentially smaller than the other one. For example, the $d$-dimensional cube $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{\infty}\leq 1\}$ has $2^{d}$ vertices but only $2d$ facets. On the other hand, the $d$-dimensional cross-polytope $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{1}\leq 1\}$ has only $2d$ vertices, but $2^{d}$ facets. Standard references for polytope theory are Grünbaum (2003); Ziegler (2012).
Extended Formulations.
Sometimes, a more compact way of representing a polyhedron $P$ can be obtained by representing it as the projection of a higher-dimensional polyhedron $Q\subseteq\mathbb{R}^{e}$ with $e\geq d$. Such a polytope $Q$ with $\pi(Q)=P$ for an affine projection $\pi$ is called an extended formulation of $P\subseteq\mathbb{R}^{d}$. The extension complexity of $P$ is defined as $\operatorname{xc}(P)=\min\{h(Q)\mid\pi(Q)=P\}$. The extension complexity is upper bounded by $\min\{v(P),h(P)\}$, but can be exponentially smaller. For example, the spanning tree polytope of a graph with $n$ vertices has extension complexity $\mathcal{O}(n^{3})$ (Martin, 1991), but the polytope itself has exponentially many vertices and facets in $n$. See Conforti et al. (2013) for a (not very recent) survey on extended formulations.
Support Functions.
@@
The inverse map that maps a support function $f_{P}$ to its unique associated polytope $P$ is well-studied in tropical geometry. In fact, borrowing the name from (tropical) polynomials, we call $P$ the Newton polytope of $f_{P}$. The connection between CPWL functions and their Newton polytopes has previously been used in the study of neural networks and their expressivity (Zhang et al., 2018; Montúfar et al., 2022; Haase et al., 2023; Hertrich et al., 2023; Brandenburg et al., 2024b; Misiakos et al., 2022).
Neural Networks.
For an introduction to neural networks, we refer to Goodfellow et al. (2016). In this paper, we focus on so-called rank-$k$ maxout neural networks for some natural number $k\geq 2$. They generalize the well-known neural networks with rectified linear unit (ReLU) activations. Such a neural network is based on a directed, acyclic graph with node set $V$ and arcs $A\subseteq V\times V$, where the nodes are called neurons and act as computational units. The $d\geq 1$ neurons with in-degree zero are called input neurons, all other $s\geq 1$ neurons are maxout units. We assume that among the maxout units there exists exactly one neuron with out-degree zero, the output neuron. All neurons that are neither input nor output neurons are called hidden neurons. The size of the neural network is defined as the number $s$ of maxout units.
Each neuron $v$ in the neural network defines a function $z_{v}\colon\mathbb{R}^{d}\to\mathbb{R}$ as follows. Each input neuron $v$ is associated with one element $x_{i}$ of the input vector $x\in\mathbb{R}^{d}$ and simply outputs $z_{v}(x)=x_{i}$. Each maxout unit $v$ comes with a tuple of weights $w_{uv}^{i}\in\mathbb{R}$ for $i=1,\dots,k$ and all $u\in\delta^{\mathrm{in}}_{v}$, where $\delta^{\mathrm{in}}_{v}$ is the set of in-neighbors of $v$. The maxout neuron $v$ represents the following expression dependent on the outputs of its in-neighbors:
@@
Finally, the output of the network is defined to be the output $z_{v}$ for the output neuron $v$. Note that our neural network represents a scalar-valued function as we assumed that there is only a single output neuron. A maxout network is called monotone if all weights of all maxout neurons are nonnegative (Mikulincer and Reichman, 2022). It is called an input-convex neural network (ICNN) if negative weights are only allowed on arcs leaving an input neuron (Amos et al., 2017).
For a polytope $P$, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum size of a rank-$2$ maxout network representing $f_{P}$. Moreover, we define the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum size of a rank-$2$ maxout ICNN representing $f_{P}$. We will see that, if $f_{P}$ is a monotone function, this exactly equals the minimum size of a monotone rank-$2$ maxout network. While the previous use of the word ‘monotone’ was for neural networks, recall that a multivariate function is called monotone if, for $x,y\in\mathbb{R}^{d}$, the componentwise inequality $x\leq y$ implies $f(x)\leq f(y)$.
Remark 2.1 (biases).
In practice, the units usually also involve biases. In the following, we will mainly focus on support functions of polytopes, which are positively homogeneous. Therefore, we can omit the bias in the definition of neural networks without loss of generality, compare Hertrich et al. (2023, Proposition 2.3). With our definition of a maxout network, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ can be represented by such a network if and only if it is CPWL and positively homogeneous.
Remark 2.2 (maxout ranks and ReLU).
@@
3. Lower Bounds for Monotone and Input-Convex Neural Networks
The goal of this section is to prove the following theorem and use it to derive exponential lower bounds on monotone and input-convex neural networks. Here and in the following, we defer some of the proofs to Appendix A.